# Bidirectional Hierarchical Protein Multi-Modal Representation Learning

### v1

@@ 1 Introduction
Proteins are essential building blocks of life. While proteins can be represented as one-dimensional sequential data, their complex three-dimensional structures and dynamic nature underscore their vast functional diversity. A thorough understanding of protein 3D structures is critical for unraveling disease mechanisms and advancing drug discovery. Consequently, extensive research has been conducted on protein 3D structure representation learning, demonstrating its effectiveness across diverse protein analysis tasks [2, 19, 23]. With advancements in deep learning, 3D geometric graph neural networks (GGNNs) have been developed to model protein structural information, yielding significant improvements in prediction tasks involving proteins [9, 25, 19, 21]. However, the limited availability of labeled data constrains the power of GGNNs. In addition, existing GGNNs are proven to be unaware of the positional order within the protein sequence [21].
@@
On the other hand, protein folding models [11, 13], which predict 3D structures from protein sequences, highlight the rich information embedded in one-dimensional sequential data. Inspired by the success of large pretrained language models (LLMs) in natural language processing [14, 15], researchers have adapted LLMs for protein representation learning using protein sequences. These protein language models (pLMs) treat protein sequences as a language, with individual amino acids as tokens. Prominent advancements in this area include UniRep [1], ProtTrans [8], and ESM [13]. Although pLMs benefit from the versatile Transformer architecture and the relatively greater availability of unlabeled sequence data, a model based on sequence prediction alone lacks the structural information and hierarchical representation of proteins, which restricts the utilization of labeled structural data and may lack the inductive bias to represent proteins consistent with physical and chemical constraints. Thus, adapting pLMs to tasks involving structural input, such as protein structure and protein-protein interaction prediction remains challenging.
Therefore, integrating diverse modalities of data representation offers a promising avenue to enrich protein analysis. The serial fusion framework proposed in previous work [21, 26, 24] is one way to combine representations of pLMs and GNNs for supervised learning applications, as depicted in Fig. 1(A). Initially, protein sequences are processed through a pLMs to generate detailed per-residue representations. These representations are then utilized as node features in 3D protein graphs, which are further analyzed using a GNN. This integration ensures that the information captured by pLMs enrich the structural analysis performed by GNN, potentially leading to more accurate and insightful predictions.
A notable example of serial fusion is ESM-GearNet model [26]. This model incorporates the output of the ESM into GearNet [25], by substituting GearNet’s node features with those derived from ESM. The resultant representation benefits from the deep evolutionary insights encoded by pLMs, demonstrating the potential of combining pLMs with GNNs for advanced protein representation. However, the reliance of ESM-GearNet on self-supervised learning for pre-training poses questions about its adaptability and efficacy in supervised learning contexts. Furthermore, there are two inadequacies of the serial fusion method as shown in Fig. 1(A). One is that pLMs do not receive the structural information from GNN, Thus the interaction between two branches is only unidirectional. The second is that the exchange of information between two branches happens only once, which may limit how much the system can fully benefit from different but complementary views of the same object at different hierarchical layer.
@@
Local Bi-Hierarchical Fusion with gating. To effectively combine features from these representations, we utilize a gating mechanism, as shown in Figure 2(A). This method dynamically adjusts the weight of each modality’s contribution, improving the integration process and clarifying the influence of each modality on the final prediction. The gated fusion layer modifies the traditional addition-based fusion approach by controlling the information flow through gates, drawing inspiration from RPVNet [22]. This technique not only enables seamless integration of sequence and graph data but also enhances the model’s interpretability by highlighting the contributions of different modalities on the final outcomes. See §2.3.1 for more details.
@@
We evaluate our methodologies across a diverse array of benchmarks, including model quality assessment, protein-ligand binding affinity, reaction prediction, protein-protein binding site prediction, and B cell epitopes (BCE) prediction. Our comprehensive experiments demonstrate that our Bi-Hierarchical Fusion approach, which including token-wise and global information exchange, surpasses the previous state-of-the-art in protein representation learning, Pronet, and serial fusion across various tasks that require both structural and sequential knowledge. This underscores that our Bi-Hierarchical Fusion technique facilitates more effective knowledge exchange between different branches of protein representation compared to serial fusion approach. Our findings contribute to a better understanding of how to effectively utilize burgeoning geometric deep learning, well-established protein language models, and enhance the integration of various protein modalities.
@@ 2.1 Background
Notations. To model protein representations with 3D structures, we represent a protein as a 3D graph $G=(\mathcal{V},\mathcal{E},\mathcal{P})$. Here, $\mathcal{V}=\left\{{\mathbf{v}_{i}}\right\}_{i=1,...,n}$ represents the set of node features, with each $\mathbf{v}_{i}\in\mathbb{R}^{d_{v}}$ indicating the feature vector for node $i$. =$\mathcal{E}=\left\{e_{ij}\right\}_{i,j=1,\cdots,n}$ comprises the set of edge features, where $e_{ij}\in\mathbb{R}$ corresponds to the feature vector for edge $ij$. $\mathcal{P}=\left\{P_{i}\right\}_{i=1,\cdots,n}$ denotes the set of position matrices, with each $P_{i}\in\mathbb{R}^{k_{i}\times 3}$ representing the position matrix for node $i$. The value of $k_{i}$ varies across different applications. For instance, in the context of molecules where each atom is considered a node, $k_{i}$ is 1 for each node $i$. Conversely, in proteins where each amino acid is treated as a node, $k_{i}$ corresponds to the number of atoms in amino acid $i$.
In this graph, each amino acid is a node, and edges are established between nodes if the distance between them is less than a certain cutoff radius. Each node $i$ has a feature $\mathbf{v}i$, which is a one-hot encoding of its amino acid type. Each edge $ij$ has a feature $\mathbf{e}_{ij}$, representing the embedding of the sequential distance $j-i$, consistent with prior research. Additionally, the position matrix $P_{i}$ for a node $i$ contains the coordinates for all atoms of the amino acid when available, arranged in a predefined atom order. For instance, in the amino acid alanine, the sequence of atoms in the position matrix is $N,C_{\alpha},C,O$, and $C_{\beta}$.
Complete Geometric Representations. As described in [18], a geometric transformation $\mathcal{F}\left({\cdot}\right)$ is considered complete if it holds that for any two 3D graphs $G^{1}=\left({\mathcal{V},\mathcal{E},\mathcal{P}^{1}}\right)$ and $G^{2}=\left({\mathcal{V},\mathcal{E},\mathcal{P}^{2}}\right)$, the condition $\mathcal{F}\left({G^{1}}\right)=F\left({G^{2}}\right)\Leftrightarrow\exists R\in SE\left({3}\right)$ implies there exists a transformation $R\in SE\left({3}\right)$ such that for every $i$ from 1 to $n$, $P_{i}^{1}=R\left({P^{2}_{i}}\right)$. The group $SE\left({3}\right)$ represents the Sepecial Euclidean group that accounts for all possible rotations and translations in three dimensions. A complete geometric representation inherently possesses rotation and translation invariance, reflecting the natural properties of proteins and offering a robust framework for analyzing protein structures.
@@ 2.2 Leveraging Protein Large Language Models and Graph Neural Networks
For our framework, we select ESM [13] as the transformer-based protein language model, ProNet [19] for protein GNN. ESM is chosen for its proven excellence in tasks such as protein structure prediction with ESMFold [13] and protein function prediction in Gearnet-ESM [26]. Its capability to encode complex biological data into meaningful representations positions it as an ideal candidate for enhancing protein representation learning and, consequently, benefitting a variety of protein representation learning tasks, such as binding affinity predictions. ProNet stands out for its geometric representation capabilities at the amino acid level, employing a comprehensive message passing scheme to achieve a full 3D protein graph representation. This aspect is critical for our framework, as ProNet’s geometric transformation $F(\cdot)$ ensures SE(3) invariance, essential for maintaining the accuracy of protein representations despite rotations and translations [6]. This invariance is crucial for accurately comparing protein structures by eliminating discrepancies caused by orientation or positional differences.
@@ 3 Results and discussion
RMSE $\downarrow$ $R_{p}$uparrow $R_{s}$uparrow RMSE $\downarrow$ $R_{p}$uparrow $R_{s}$uparrow
@@
Table 1: Results on protein-ligand binding affinity prediction task. For baselines, We took the results from the paper of ProNet [19]. The top rows all use ProNet-Amino Acid as the GNN for proteins, and the bottom rows all use ProNet-All Atom. All our fusion models use ESM-2 for pLMs. Bolded numbers are the best performance within the comparison group. The top two results are highlighted as 1st and 2nd.
3.1 Experimental setup
We evaluate our proposed framework across five established protein benchmarks, including Model Quality Assessment (MQA), Protein-Ligand Binding Affinity (LBA), Reaction Prediction, Protein-Protein Binding Site Prediction (PPBS), and B-cell Epitope (BCE) Prediction (We defer the detail of datasets, tasks and results in §3.2). In these evaluations, we employ ProNet [19] for its GNN architectures and ESM-2 [13] as the pretrained protein language models (pLMs). ProNet is noted for its hierarchical protein representations and an extensive message passing system that captures a complete 3D protein graph representation. A vital component of our framework involves ProNet’s geometric transformation $F(\cdot)$, which guarantees SE(3) invariance. This invariance is critical for preserving the accuracy of protein representations regardless of their rotations and translations [6]. Ensuring this invariance is crucial for consistent and accurate comparisons of protein structures by eliminating variability due to orientation or positional differences.
Using a pretrained GNN and pLMs, we conduct experiments with serial fusion [21], where representations from the pLMs are fed into the GNN as features. Both ProNet and Serial Fusion achieved previous state-of-the-art performance across various tasks and served as strong competitive baselines. Additionally, we explore our local Bi-Hierarchical Fusion with gating and global Bi-Hierarchical Fusion employing multi-head self-attention. Furthermore, for the ligand binding affinity task, we utilize the same GNN for the ligand as described in the studies we reference  [19, 16].
3.2 Tasks, datasets and experiment results
3.2.1 Single-protein representation task
Reaction Classification. Reaction classification prediction involves determining the specific biochemical reaction catalyzed by an enzyme, a task critical for understanding metabolic pathways and designing enzyme-targeted drugs. Enzymes, which serve as biological catalysts, are categorized by enzyme commission (EC) numbers based on the reactions they facilitate. We utilize the dataset and experimental setup from  Hermosilla et al. [10] to evaluate our methods. This dataset comprises 37,428 proteins across 384 EC numbers [3, 5]. For reaction prediction, we assess performance using accuracy metrics.
Method pLMs Accuracy
ProNet (Amino Acid) $\times$ 0.79
Serial Fusion $\checkmark$ 0.8105
Local Bi-Hierarchical Fusion $\checkmark$ 0.8757
Global Bi-Hierarchical Fusion $\checkmark$ 0.8412
Table 2: Comparison of methods on reaction prediction. GNN uses ProNet (Amino Acid) as the baseline backbone. The top two results are highlighed as 1st and 2nd.
Model Quality Assessment. Model Quality Assessment (MQA) plays a critical role in structure prediction by selecting the best structural model of a protein from a large pool of candidate structures [4]. For many recently solved but unreleased protein structures, structure generation algorithms produce an extensive set of candidate models. MQA methods are evaluated based on their ability to predict the Global Distance Test Total Score (GDT-TS) of a candidate structure relative to the experimentally determined structure of the target protein. The evaluation of MQA approaches often relies on databases comprising all structural models submitted to the Critical Assessment of Structure Prediction (CASP) [12] experiments.
Method pLMs MSE $\downarrow$ $R_{p}$uparrow $R_{s}$uparrow
ProNet (Amino Acid) $\times$ 0.1934 0.5479 0.5948
Serial fusion $\checkmark$ 0.1915 0.6024 0.5942
Local Bi-Hierarchical Fusion $\checkmark$ 0.1846 0.6075 0.5969
Global Bi-Hierarchical Fusion $\checkmark$ 0.1881 0.5977 0.5961
Table 3: Comparison of methods on model quality assessment (MQA). GNN uses ProNet (Amino Acid) as the baseline backbone. The top two results are highlighted as 1st and 2nd.

### v2

@@ 1 Introduction
Proteins are essential building blocks of life. While proteins can be represented as one-dimensional sequential data, their complex three-dimensional structures and dynamic nature underscore their vast functional diversity. A thorough understanding of protein 3D structures is critical for unraveling disease mechanisms and advancing drug discovery. Consequently, extensive research has been conducted on protein 3D structure representation learning, demonstrating its effectiveness across diverse protein analysis tasks (Baldassarre et al., 2021; Wang et al., 2023; Yang et al., 2023). With advancements in deep learning, 3D geometric graph neural networks (GGNNs) have been developed to model protein structural information, yielding significant improvements in prediction tasks involving proteins (Fan et al., 2022; Zhang et al., 2022; Wang et al., 2023; Wu et al., 2023). However, the limited availability of labeled data constrains the power of GGNNs. In addition, existing GGNNs are proven to be unaware of the positional order within the protein sequence (Wu et al., 2023).
@@
On the other hand, protein folding models (Jumper et al., 2021; Lin et al., 2023), which predict 3D structures from protein sequences, highlight the rich information embedded in one-dimensional sequential data. Inspired by the success of large pretrained language models (LLMs) in natural language processing (Radford et al., 2019; Raffel et al., 2020), researchers have adapted LLMs for protein representation learning using protein sequences. These protein language models (pLMs) treat protein sequences as a language, with individual amino acids as tokens. Prominent advancements in this area include UniRep (Alley et al., 2019), ProtTrans (Elnaggar et al., 2021), and ESM (Lin et al., 2023). Although pLMs benefit from the versatile Transformer architecture and the relatively greater availability of unlabeled sequence data, a model based on sequence prediction alone lacks the structural information and hierarchical representation of proteins, which restricts the utilization of labeled structural data and may lack the inductive bias to represent proteins consistent with physical and chemical constraints. Thus, adapting pLMs to tasks involving structural input, such as protein structure and protein-protein interaction prediction remains challenging.
Therefore, integrating diverse modalities of data representation offers a promising avenue to enrich protein analysis. The serial fusion framework proposed in previous work (Wu et al., 2023; Zhang et al., 2023b, a) is one way to combine representations of pLMs and GNNs for supervised learning applications, as depicted in Fig. 1(A). Initially, protein sequences are processed through a pLMs to generate detailed per-residue representations. These representations are then utilized as node features in 3D protein graphs, which are further analyzed using a GNN. This integration ensures that the information captured by pLMs enrich the structural analysis performed by GNN, potentially leading to more accurate and insightful predictions.
A notable example of serial fusion is ESM-GearNet model (Zhang et al., 2023b). This model incorporates the output of the ESM into GearNet (Zhang et al., 2022), by substituting GearNet’s node features with those derived from ESM. The resultant representation benefits from the deep evolutionary insights encoded by pLMs, demonstrating the potential of combining pLMs with GNNs for advanced protein representation. However, the reliance of ESM-GearNet on self-supervised learning for pre-training poses questions about its adaptability and efficacy in supervised learning contexts. Furthermore, there are two inadequacies of the serial fusion method as shown in Fig. 1(A). One is that pLMs do not receive the structural information from GNN, Thus the interaction between two branches is only unidirectional. The second is that the exchange of information between two branches happens only once, which may limit how much the system can fully benefit from different but complementary views of the same object at different hierarchical layer.
@@
Local Bi-Hierarchical Fusion with gating. To effectively combine features from these representations, we utilize a gating mechanism, as shown in Figure 2(A). This method dynamically adjusts the weight of each modality’s contribution, improving the integration process and clarifying the influence of each modality on the final prediction. The gated fusion layer modifies the traditional addition-based fusion approach by controlling the information flow through gates, drawing inspiration from RPVNet (Xu et al., 2021). This technique not only enables seamless integration of sequence and graph data but also enhances the model’s interpretability by highlighting the contributions of different modalities on the final outcomes. See §2.3.1 for more details.
@@
Generalizable Insights about Machine Learning in the Context of Healthcare
Enhancing protein representation has direct impact on drug discovery, structural biology, and immunotherapy, and plays a crucial role in advancing AI applications in clinical research and practice, including drug and vaccine development. In this work, we introduce the Bi-Hierarchical Fusion framework for protein representation learning. Building on this framework, we develop two variants: local Bi-Hierarchical Fusion with gating mechanisms and global Bi-Hierarchical Fusion with multi-head attention. We evaluate our methodologies across a diverse array of benchmarks, including model quality assessment, protein-ligand binding affinity, reaction prediction, protein-protein binding site prediction, and B cell epitopes (BCE) prediction. Our comprehensive experiments demonstrate that our Bi-Hierarchical Fusion approach, which includes token-wise and global information exchange, surpasses the previous state-of-the-art in protein representation learning, Pronet, and serial fusion across various tasks that require both structural and sequential knowledge. This underscores that our Bi-Hierarchical Fusion technique facilitates more effective knowledge exchange between different branches of protein representation compared to the serial fusion approach. Our findings contribute to a better understanding of how to effectively utilize burgeoning geometric deep learning, well-established protein language models, and enhance the integration of various protein modalities.
@@ 2.1 Background
Notations. To model protein representations with 3D structures, we represent a protein as a 3D graph $G=(\mathcal{V},\mathcal{E},\mathcal{P})$. Here, $\mathcal{V}=\left\{{\mathbf{v}_{i}}\right\}_{i=1,...,n}$ represents the set of node features, with each $\mathbf{v}_{i}\in\mathbb{R}^{d_{v}}$ indicating the feature vector for node $i$. $\mathcal{E}=\left\{e_{ij}\right\}_{i,j=1,\cdots,n}$ comprises the set of edge features, where $e_{ij}\in\mathbb{R}$ corresponds to the feature vector for edge $ij$. $\mathcal{P}=\left\{P_{i}\right\}_{i=1,\cdots,n}$ denotes the set of position matrices, with each $P_{i}\in\mathbb{R}^{k_{i}\times 3}$ representing the position matrix for node $i$. The value of $k_{i}$ varies across different applications. For instance, in the context of molecules where each atom is considered a node, $k_{i}$ is 1 for each node $i$. Conversely, in proteins where each amino acid is treated as a node, $k_{i}$ corresponds to the number of atoms in amino acid $i$.
In this graph, each amino acid is a node, and edges are established between nodes if the distance between them is less than a certain cutoff radius. Each node $i$ has a feature $\mathbf{v}i$, which is a one-hot encoding of its amino acid type. Each edge $ij$ has a feature $\mathbf{e}_{ij}$, representing the embedding of the sequential distance jij-i, consistent with prior research. Additionally, the position matrix $P_{i}$ for a node $i$ contains the coordinates for all atoms of the amino acid when available, arranged in a predefined atom order. For instance, in the amino acid alanine, the sequence of atoms in the position matrix is $N,C_{\alpha},C,O$, and $C_{\beta}$.
Complete Geometric Representations. As described in (Wang et al., 2022), a geometric transformation $\mathcal{F}\left({\cdot}\right)$ is considered complete if it holds that for any two 3D graphs $G^{1}=\left({\mathcal{V},\mathcal{E},\mathcal{P}^{1}}\right)$ and $G^{2}=\left({\mathcal{V},\mathcal{E},\mathcal{P}^{2}}\right)$, the condition $\mathcal{F}\left({G^{1}}\right)=F\left({G^{2}}\right)\Leftrightarrow\exists R\in SE\left({3}\right)$ implies there exists a transformation $R\in SE\left({3}\right)$ such that for every $i$ from 1 to $n$, $P_{i}^{1}=R\left({P^{2}_{i}}\right)$. The group $SE\left({3}\right)$ represents the Sepecial Euclidean group that accounts for all possible rotations and translations in three dimensions. A complete geometric representation inherently possesses rotation and translation invariance, reflecting the natural properties of proteins and offering a robust framework for analyzing protein structures.
@@ 2.2 Leveraging Protein Large Language Models and Graph Neural Networks
For our framework, we select ESM (Lin et al., 2023) as the transformer-based protein language model, ProNet (Wang et al., 2023) for protein GNN. ESM is chosen for its proven excellence in tasks such as protein structure prediction with ESMFold (Lin et al., 2023) and protein function prediction in Gearnet-ESM (Zhang et al., 2023b). Its capability to encode complex biological data into meaningful representations positions it as an ideal candidate for enhancing protein representation learning and, consequently, benefitting a variety of protein representation learning tasks, such as binding affinity predictions. ProNet stands out for its geometric representation capabilities at the amino acid level, employing a comprehensive message passing scheme to achieve a full 3D protein graph representation. This aspect is critical for our framework, as ProNet’s geometric transformation $F(\cdot)$ ensures SE(3) invariance, essential for maintaining the accuracy of protein representations despite rotations and translations (Defresne et al., 2021). This invariance is crucial for accurately comparing protein structures by eliminating discrepancies caused by orientation or positional differences.
@@ 3 Results and discussion
3.1 Experimental setup
We evaluate our proposed framework across five established protein benchmarks, including Model Quality Assessment (MQA), Protein-Ligand Binding Affinity (LBA), Reaction Prediction, Protein-Protein Binding Site Prediction (PPBS), and B-cell Epitope (BCE) Prediction (We defer the detail of datasets, tasks and results in §3.2). In these evaluations, we employ ProNet (Wang et al., 2023) for its GNN architectures and ESM-2 (Lin et al., 2023) as the pretrained protein language models (pLMs). ProNet is noted for its hierarchical protein representations and an extensive message passing system that captures a complete 3D protein graph representation. A vital component of our framework involves ProNet’s geometric transformation $F(\cdot)$, which guarantees SE(3) invariance. This invariance is critical for preserving the accuracy of protein representations regardless of their rotations and translations (Defresne et al., 2021). Ensuring this invariance is crucial for consistent and accurate comparisons of protein structures by eliminating variability due to orientation or positional differences.
Using a pretrained GNN and pLMs, we conduct experiments with serial fusion (Wu et al., 2023), where representations from the pLMs are fed into the GNN as features. Both ProNet and Serial Fusion achieved previous state-of-the-art performance across various tasks and served as strong competitive baselines. Additionally, we explore our local Bi-Hierarchical Fusion with gating and global Bi-Hierarchical Fusion employing multi-head self-attention. Furthermore, for the ligand binding affinity task, we utilize the same GNN for the ligand as described in the studies we reference  (Wang et al., 2023; Somnath et al., 2021).
3.2 Tasks, datasets and experiment results
3.2.1 Single-protein representation task
Method pLMs Accuracy
ProNet (Amino Acid) $\times$ 0.79
Serial Fusion $\checkmark$ 0.8105
Local Bi-Hierarchical Fusion $\checkmark$ 0.8757
Global Bi-Hierarchical Fusion $\checkmark$ 0.8412
Table 1: Comparison of methods on reaction prediction. GNN uses ProNet (Amino Acid) as the baseline backbone. The top two results are highlighted as 1st and 2nd.
Reaction Classification. Reaction classification prediction involves determining the specific biochemical reaction catalyzed by an enzyme, a task critical for understanding metabolic pathways and designing enzyme-targeted drugs. Enzymes, which serve as biological catalysts, are categorized by enzyme commission (EC) numbers based on the reactions they facilitate. We utilize the dataset and experimental setup from  Hermosilla et al. (2020) to evaluate our methods. This dataset comprises 37,428 proteins across 384 EC numbers (Berman et al., 2002; Dana et al., 2019). For reaction prediction, we assess performance using accuracy metrics.
Method pLMs MSE $\downarrow$ $R_{p}$ $\uparrow$ $R_{s}$ $\uparrow$
ProNet (Amino Acid) $\times$ 0.1934 $\pm$ 0.006 0.5479 $\pm$ 0.034 0.5948 $\pm$ 0.052
Serial fusion $\checkmark$ 0.1915 $\pm$ 0.001 0.6024 $\pm$ 0.005 0.5942 $\pm$ 0.001
Local Bi-Hierarchical Fusion $\checkmark$ 0.1846 $\pm$ 0.001 0.6075 $\pm$ 0.006 0.5969 $\pm$ 0.002
Global Bi-Hierarchical Fusion $\checkmark$ 0.1881 $\pm$ 0.001 0.5977 $\pm$ 0.003 0.5961 $\pm$ 0.004
Table 2: Comparison of methods on model quality assessment (MQA). GNN uses ProNet (Amino Acid) as the baseline backbone. The top two results are highlighted as 1st and 2nd. Results on over 5 independent runs.
Model Quality Assessment. Model Quality Assessment (MQA) plays a critical role in structure prediction by selecting the best structural model of a protein from a large pool of candidate structures (Cheng et al., 2019). For many recently solved but unreleased protein structures, structure generation algorithms produce an extensive set of candidate models. MQA methods are evaluated based on their ability to predict the Global Distance Test Total Score (GDT-TS) of a candidate structure relative to the experimentally determined structure of the target protein. The evaluation of MQA approaches often relies on databases comprising all structural models submitted to the Critical Assessment of Structure Prediction (CASP) (Kryshtafovych et al., 2019) experiments.
@@ 3.2.1 Single-protein representation task
RMSE $\downarrow$ $R_{p}$ $\uparrow$ $R_{s}$ $\uparrow$ RMSE $\downarrow$ $R_{p}$ $\uparrow$ $R_{s}$ $\uparrow$
@@
Table 3: Results on protein-ligand binding affinity prediction task. For baselines, We took the results from the paper of ProNet (Wang et al., 2023). The top rows all use ProNet-Amino Acid as the GNN for proteins, and the bottom rows all use ProNet-All Atom. All our fusion models use ESM-2 for pLMs. Bolded numbers are the best performance within the comparison group. The top two results are highlighted as 1st and 2nd.
3.2.2 Protein-molecules representation task
Ligand Binding affinity. Protein-ligand binding affinity (LBA) prediction (Liu et al., 2024) is a critical task in drug discovery, as it estimates the interaction strength between a candidate drug molecule and a target protein. For this study, we utilize the PDBbind database (Wang et al., 2005), a curated resource of protein-ligand complexes sourced from the Protein Data Bank (PDB), annotated with their respective binding strengths. To ensure a robust evaluation, the dataset is partitioned such that proteins in the test set share no more than 30% or 60% sequence identity with any protein in the training set.
The advancement of interaction-free methods as illustrated in  Fig. 3 for binding affinity prediction emphasizes the need for sophisticated computational techniques independent of physical interaction data. These methods employ separate neural networks and representations for proteins and molecules, with proteins providing a more complex computational canvas due to their intricate structures and dynamic functions. In our study, we represent proteins using ProNet and ligands as outlined in (Wang et al., 2023; Somnath et al., 2021). We evaluate these methods at both the amino-acid and all-atom levels for a comprehensive assessment of binding affinity prediction. Our evaluations involve comparing root mean squared error (RMSE), Pearson correlation ($R_{p}$), and Spearman correlation ($R_{s}$), with the main results detailed in Table 3. ProNet, previously the state-of-the-art, serves as our baseline. Serial fusion improves upon this, reducing RMSE by 0.03-0.05. Our Bi-Hierarchical Fusion approaches further decrease RMSE by 0.01-0.08, with global Bi-Hierarchical Fusion achieving the lowest RMSE in the challenging 30% sequence identity split, demonstrating strong generalizability.
3.2.3 Protein-protein representation task